OpenStudy (anonymous):

PLEASE help! Probability! - A shoebox holds same-size disks. There are 5 red, 6 white, and 7 blue disks. You pick out a disk, record its color, and return it to the box. If you repeat this process 250 times, how many times can you expect to pick either a red or white disk?

OpenStudy (anonymous):

please help me!!

OpenStudy (lstrasmann):

5 + 6 + 7 = 18 P = 250([6 + 7] / 18) P = 250(13 / 18) That's all I remember sorry :)

OpenStudy (lstrasmann):

I think you have to distribute but I think distribution is for addition and subtraction only, not sure though.